OpenStudy (anonymous):

Find the domain of the composite function . Please show all of your work. f(x) = x/x+9; g(x)= 45/x+5 Find the domain of the composite function . Please show all of your work. f(x) = x/x+9; g(x)= 45/x+5 @Mathematics

OpenStudy (alfie):

\[f(g(x)) = \frac{\frac{45}{x+5}}{\frac{45}{x+5} + 9}\] LCD at the denominator to get: \[f(g(x)) = \frac{\frac{45}{x+5}}{\frac{90 + 9x}{x+5}}\] Which is... \[\frac{45}{\cancel {x+5}}\frac{\cancel {x+5}}{90+9x} \rightarrow \frac{45}{9(10+x)} \rightarrow \frac{5}{10+x}\] So.. \[10 + x \neq 0 \rightarrow x \neq -10\]

OpenStudy (anonymous):

wow, im impressed. is this the answer?

OpenStudy (alfie):

The answer of the domain is: \[x \in R \neq {(-10)}\]

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

How did you get this?!

OpenStudy (anonymous):

IM trying to figure out what the second function says. I can't read the bottom numbers. Can you tell me what they are?

OpenStudy (alfie):

f(g(x)) = \[\huge \frac{\frac{45}{x+5}}{\frac{90 + 9x}{x+5}}\] \[\frac{45}{\cancel {x+5}}\frac{\cancel {x+5}}{90+9x} \rightarrow \frac{45}{9(10+x)} \rightarrow \frac{5}{10+x}\]

OpenStudy (anonymous):

great. lol- thank you again! It was in a small print and I couldnt really tell what it was. :)

OpenStudy (alfie):

And if you have.. f(x) = x+1 and g(x) = 3... The f(g(x)) is just sticking the value of g(x) inside of the f(x) so... f(x) = 3+1 => 4. So it's just a constant function y = 4. Bye :)

OpenStudy (anonymous):

ok- i forgot to put in the original equation that is was asking for fog...will that change this any?

OpenStudy (alfie):

(f ∘ g)(x) = f(g(x)) So no problem ;)

OpenStudy (anonymous):

oh gotcha...ok- so really I dont have to worry about anything changing? lol

OpenStudy (alfie):

nope, different notation, same stuff.

OpenStudy (anonymous):

cool.

jhonyy9 (jhonyy9):

but firstly you need to suppose that x not can being equal -9 and -5 so because than those freactions are undefined

jhonyy9 (jhonyy9):

fractions

jhonyy9 (jhonyy9):

sorry